What is the eccentricity of the hyperbola given by the equation (x^2/64)-(y^2/16)=1?
Answer by lwsshak3(11628)

What is the eccentricity of the hyperbola given by the equation
(x^2/64)-(y^2/16)=1?
This is an equation of a hyperbola with horizontal transverse axis.
Its standard form: (x-h)^2/a^2-(y-k)^2/b^2=1
For given equation:
a^2=64
a=√64=8
b^2=16
b=4
c^2=a^2+b^2=64+16=80
c=√80
..
eccentricity=c/a=√80/8=4√5/8=√5/2